OpenStudy (anonymous):

6. Write the equation of a line in slope intercept form that passes through the points (-2, 7) and (3, 12)

OpenStudy (anonymous):

@03225186213

OpenStudy (anonymous):

oh this title says parralell n perpendicular lines

OpenStudy (anonymous):

but thts the problem for #6

OpenStudy (anonymous):

y-7/12-7=x-(-2)/3+2 y-7/5=x+2/5 y-7=x+2 y=x+9

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

m=y2-y1/x2-x1 for parallel lines slopes are same for perpendicular line it is negative inverse